# The role of three-dimensional effects on ion injection and acceleration in perpendicular shocks

### v1

@@ 1 Introduction
Here we perform 2D and 3D hybrid perpendicular shock simulations with a specific focus on how particles are injected, and why 3D is necessary in order to achieve efficient injection and acceleration. The paper is structured as follows: in Section 2, we describe the simulation setup; in Section 3, we examine how our results depend on the adopted resolution and discuss the physics of ion injection in perpendicular shocks; finally, in Section 4, we present our conclusions.
2 Simulation setup
@@
In contrast to PIC simulations, the hybrid approach does not resolve the electron plasma scales. Therefore, it is possible to simulate larger systems without losing important information about the shock dynamics, which is mostly controlled by the ions. The main disadvantage is that, in the absence of constraints imposed by electron-scale physics, it is not straightforward to determine the resolution required to accurately capture ion-scale dynamics. We will discuss the appropriate resolution to adopt in Section 3.1, and explain why, once the ion skin depth is properly resolved, smaller-scale structures become irrelevant for capturing ion-scale physics.
@@
Run $M_{A}$ $\Delta t$[$\omega_{c}^{-1}$] $\Delta x$ [$d_{i}$] $\varepsilon(>10E_{sh})$ $\alpha$
A 30 5 ×$\times 10^{-4}$ 0.1 0.3% 6
B 60 2 ×$\times 10^{-4}$ 0.1 0.6% 5
C 100 3 ×$\times 10^{-4}$ 0.4 10% 2.5
D 100 1 ×$\times 10^{-4}$ 0.1 0.9% 4
@@
The simulations use different numbers $N$ of cells per $d_{i}$, as discussed in Section 3.1, which results in varying the spatial resolution $\Delta x$. The transverse box extends over 20 $d_{i}$ in the transverse directions $y$ and $z$, and we adopt 8 ion particles per cell. The timestep $\Delta t$ is chosen in each simulation to satisfy the Courant-Friedrichs-Lewy (CFL) condition. We verified the robustness of our results through convergence tests in number of particles per cell, box size, and temporal resolution for each spatial resolution considered. Details of the simulations are provided in Table 1. Most of our work is based on 3D simulations, but we also use 2D simulations for comparison. All 2D simulations are performed with $\mathbf{B}_{0}$ lying in the simulation plane, since we showed in Paper I that when it is out of plane, the field undergoes simple compression at the shock, as in a laminar magnetohydrodynamic (MHD) shock.
3 Results
3.1 Simulation resolution
@@
One common assumption is that resolving sub-$d_{i}$ turbulence may not be necessary, especially since the gyroradius of the thermal ions reflected at the shock and gyrating in the upstream magnetic field is on the order of $\sim M_{A}d_{i}$, and even larger for ions accelerated through SDA. The spatial resolution is particular taxing from the computational point of view in 3D, since it constrains also the time resolution: eventually the computational cost scales as $N^{4}$. The parameters of the simulations presented here are in Table 1. We consider two resolutions, $\Delta x=0.4$ and 0.1 $d_{i}$. The former corresponds to the one used in Paper I and the latter is the one necessary to achieve a visual agreement between a 2D hybrid simulation with $M_{A}=30$ and its fully kinetic counterpart performed with Tristan-MP (A. Spitkovsky, 2005), using a mass ratio $m_{i}/m_{e}=49$. This comparison, shown in Figure 1, serves to validate resolution convergence by benchmarking the hybrid approach against simulations that include the full electron kinetic physics. Figure 1 shows also the result obtained from a 2D simulation at $M_{A}=30$ with an intermediate resolution of $\Delta x=0.2$ $d_{i}$. The resolution of $\Delta x=0.1$ $d_{i}$ accurately captures the filaments produced by the ion-Weibel instability (E. S. Weibel, 1959; A. Bohdan et al., 2021; T. Jikei et al., 2024) that mediates the shock transition. These filaments are subsequently compressed and amplified by the shock, a process not properly captured at lower or intermediate resolutions.
@@
One may wonder whether further increasing the resolution would suppress particle acceleration. This is not the case, as the resolution adopted here ($\Delta x=0.1$ $d_{i}$) is calibrated to reproduce the magnetic field structure observed in full-PIC simulations, which serve as our “ground truth” benchmark.
Figure 3: Trajectories in the $x-y$ plane (top) and $x-z$ plane (bottom) of one representative particle per simulation—both 2D (dashed) and 3D (solid), at high and low resolution of $M_{A}=30$ shock—that exhibit similar behavior in the initial stages of their evolution (see the text for details).
3.2 The life of an accelerated particle in a perpendicular shock
Figure 4: The average penetration depths of ions with energy $E=7.5E_{sh}$ are shown for $M_{A}=30$ (red lines) and 100 (black lines) in units of $r_{L}^{*}=r_{L}(7.5E_{sh})$ (see the text for details). These represent the distances traveled downstream where $p_{x}$ changes sign. They are reported for 2D and 3D simulations at low and high resolutions. The hatched bands mark the $2\sigma$ intervals of the penetration distance distributions. The dotted line represents the 2D return horizon, as defined in Section 3.3.
To investigate the differences in particle injection between 2D and 3D simulations at both low and high resolutions, we present in Figure 3 the $x$$y$ (top) and $x–z$ (bottom) trajectories of one representative particle from each simulation,comparing 2D and 3D, both at high and low resolution for $M_{A}=30$. These particles are selected because they exhibit similar early-time behavior, including comparable distances traveled ahead of the shock after the first reflection, similar energies after the first gyration, low $y$-momentum following the initial cycle, and similar maximum penetration depths into the downstream region. These “doppelgänger” particles serve as analogs across different simulation setups, enabling a direct comparison of their dynamics.
@@ 3.2 The life of an accelerated particle in a perpendicular shock
Each particle undergoes a first reflection at the shock and propagates a similar distance upstream, between $17$ and $20$ $d_{i}$, gaining energy from the motional electric field. Following their initial reflection, they predominantly gyrate in the $x–z$ plane due to the background magnetic field $\mathbf{B_{0}}=B_{0}\mathbf{\hat{y}}$, with a positive momentum $p_{z}$ and a small $p_{y}$, reaching energies around $7E_{sh}$ in this first SDA cycle. After this stage their evolutions begin to diverge. Once they penetrate in the downstream, all particles travel comparable distances before their $p_{x}$ switches from negative back to positive; we call this distance, which ranges between 20 and 25 $d_{i}$ downstream of the shock, the maximum penetration depth.
@@
3.3 The return horizon
To assess this porosity argument, we perform two experiments in which test particles (40000 particles in each test) are propagated through the electromagnetic fields extracted from the simulations. We use ions with energy $E=7.5\,E_{sh}$; particles with this energy have already undergone the first SDA cycle and above this threshold differences between 2D and 3D spectra become significant.
@@ 3.3 The return horizon
We begin with the first test. Due to the upstream fields, reflected thermal particles predominantly gyrate in the $x–z$ plane and enter the shock with large $p_{x,z}$ components and a smaller $p_{y}$ after the first SDA cycle. We initialize test particles at the shock and model their momentum by sampling $p_{x}^{2}$, $p_{y}^{2}$, and $p_{z}^{2}$ from uniform distributions. The sampling intervals are based on the phase spaces measured in the simulations; all test particles are constrained to have the same energy of $E=7.5\,E_{sh}$. Specifically, we use $p_{y}^{2}/p^{2}\in[0.0\,;0.2]$ with both positive and negative $p_{y}$, $p_{x}^{2}/p^{2}\in[0.0\,;0.4]$ with $p_{x}<0$ (since the particle is entering the shock), and $p_{z}>0$ obtained as $p_{z}=\sqrt{p^{2}-p_{x}^{2}-p_{y}^{2}}$.
@@
[…]
@@
Figure 5 shows the percentage of particles with initial energy $E=7.5\,E_{sh}$ that, starting from various distances behind the shock, are able to return upstream. Results are presented for both 2D and 3D simulations at high (left panel) and low (right panel) resolution for $M_{A}=30$ and $M_{A}=100$. A clear contrast emerges: in 2D, the return probabilities rapidly drop to zero at the return horizon situated at $1.31\,r_{L}^{*}$ for both values of $M_{A}$, indicating a strong advective effect that prevents particles from escaping the downstream region. In 3D, however, a significant fraction of particles starting well beyond $1.31\,r_{L}^{*}$ are able to return upstream, consistent with the behavior observed in the “doppelgänger” particle trajectories of Section 3.2. The 2D return horizon approximately coincides with the lower boundary of the $2\sigma$ band of the distribution of penetration depths, implying that most particles propagate downstream to distances beyond the return horizon, as shown in Figure 4. This explains the inability of particles to return upstream in 2D.
Figure 6: Spatial distribution (white contours) in the $y$$z$ plane of test particles from our horizon exercise (based on the setup used for Figure 5) that successfully return upstream. The snapshot is taken at a position 0.3 $r_{L}^{*}$ behind the shock (with the test particles initially located at 1.6 $r_{L}^{*}$), for both high-resolution (left) and low-resolution (right) simulations at $M_{A}=30$. This distribution is compared to a snapshot of $B_{\perp}$ at the same location, illustrating that particles capable of returning upstream tend to avoid regions of strong $B_{\perp}$.
@@ 3.4 The shock porosity
Figure 7: Filling factor $F(x,y,z)$ distribution obtained for 2D (top row) and 3D (bottom row) at $M_{A}=30$ high (left) and low resolution (right) as a function of the downstream distance from the shock. $F(x,y,z)$ is computed at each $y$$z$ coordinate for different $x$ from the shock. From this sample, the whole distribution is obtained. A small value of $F$ implies the presence of regions of weak $B_{\perp}/B_{0}$. The comparison clearly shows that in 3D, $F$ tends to be systematically lower than in 2D, indicating enhanced porosity.
To support this argument, we consider test particles that successfully return upstream from an initial distance of 1.6 $r_{L}^{*}$; their spatial distribution in the $y$$z$ plane at $x_{sh}-x=0.3\,r_{L}^{*}$ is shown in Figure 6 as white contours, together with the slice of $B_{\perp}$ taken at the same location, for both high- and low-resolution simulations with $M_{A}=30$. It is evident, especially for the low resolution simulation, that particles capable of returning upstream tend to avoid regions of strong $B_{\perp}$, instead favoring areas where $B_{\perp}$ is small.
To quantify the occurrence of such regions, we perform a statistical analysis of the filling factor of small-$B_{\perp}$ regions in the near downstream. In particular, we assess the shock porosity by searching for extended regions—“channels” or “tubes”—at fixed $y$ and $z$ coordinates where $B_{\perp}$ remains consistently below a defined threshold along the $x$ direction, all the way to the shock. The presence of such structures implies reduced magnetic confinement and enables particle propagation back towards the shock. Although particle trajectories are fully three-dimensional and not confined to fixed $y$$z$ locations, this test provides both qualitative and quantitative evidence for the existence of preferential escape pathways.
@@
$f(x^{\prime},y,z)=\begin{cases}1&\text{if }B_{\perp}(x^{\prime},y,z)/B_{0}<T,\\ 0&\text{otherwise}.\end{cases}$
@@
In 2D simulations, the translational invariance in the $z$ direction implies that the magnetic field is organized in ”walls” that inhibit the propagation of particles back towards the upstream. Conversely, in 3D, the magnetic field lines can bend and twist in all directions, creating “tubes” with small $B_{\perp}$ in the $y$$z$ plane through which particles may escape, as illustrated in Figure 6. In 2D the field topology resembles stacked “lasagna” sheets in the $y$$z$ plane that block motion, whereas in 3D, the structure resembles porous “maccheroni” tubes (with axis along $x$) that can form extended channels with low $B_{\perp}$, enhancing porosity in the shock downstream and enabling particle injection (F. C. Jones et al., 1998). The structure of $B_{\perp}$ in the downstream region further supports this interpretation (see the Appendix).

### v2

@@ I Introduction
Here we perform 2D and 3D hybrid perpendicular shock simulations with a specific focus on how particles are injected, and why 3D is necessary in order to achieve efficient injection and acceleration. The paper is structured as follows: in Section II, we describe the simulation setup; in Section III, we examine how our results depend on the adopted resolution and discuss the physics of ion injection in perpendicular shocks; finally, in Section IV, we present our conclusions.
II Simulation setup
@@
In contrast to PIC simulations, the hybrid approach does not resolve the electron plasma scales. Therefore, it is possible to simulate larger systems without losing important information about the shock dynamics, which is mostly controlled by the ions. The main disadvantage is that, in the absence of constraints imposed by electron-scale physics, it is not straightforward to determine the resolution required to accurately capture ion-scale dynamics. We will discuss the appropriate resolution to adopt in Section III.1, and explain why, once the ion skin depth is properly resolved, smaller-scale structures become irrelevant for capturing ion-scale physics.
@@
Run $M_{A}$ $\Delta t$$\omega_{c}^{-1}$] $\Delta x$ $d_{i}$] $\varepsilon(>10E_{sh})$ $\alpha$
A 30 5 $\times 10^{-4}$ 0.1 0.3% 6
B 60 2 $\times 10^{-4}$ 0.1 0.6% 5
C 100 3 $\times 10^{-4}$ 0.4 10% 2.5
D 100 1 $\times 10^{-4}$ 0.1 0.9% 4
@@
The simulations use different numbers $N$ of cells per $d_{i}$, as discussed in Section III.1, which results in varying the spatial resolution $\Delta x$. The transverse box extends over 20 $d_{i}$ in the transverse directions $y$ and $z$, and we adopt 8 ion particles per cell. The timestep $\Delta t$ is chosen in each simulation to satisfy the Courant-Friedrichs-Lewy (CFL) condition. We verified the robustness of our results through convergence tests in number of particles per cell, box size, and temporal resolution for each spatial resolution considered. Details of the simulations are provided in Table 1. Most of our work is based on 3D simulations, but we also use 2D simulations for comparison. All 2D simulations are performed with $\mathbf{B}_{0}$ lying in the simulation plane, since we showed in Paper I that when it is out of plane, the field undergoes simple compression at the shock, as in a laminar magnetohydrodynamic (MHD) shock.
III Results
III.1 Simulation resolution
@@
One common assumption is that resolving sub-$d_{i}$ turbulence may not be necessary, especially since the gyroradius of the thermal ions reflected at the shock and gyrating in the upstream magnetic field is on the order of $\sim M_{A}d_{i}$, and even larger for ions accelerated through SDA. The spatial resolution is particular taxing from the computational point of view in 3D, since it constrains also the time resolution: eventually the computational cost scales as $N^{4}$. The parameters of the simulations presented here are in Table 1. We consider two resolutions, $\Delta x=0.4$ and 0.1 $d_{i}$. The former corresponds to the one used in Paper I and the latter is the one necessary to achieve a visual agreement between a 2D hybrid simulation with $M_{A}=30$ and its fully kinetic counterpart performed with Tristan-MP (A. Spitkovsky, 2005), using a mass ratio $m_{i}/m_{e}=49$. We have also tested a higher mass ratio, $m_{i}/m_{e}=256$, which yields results that are nearly identical at ion kinetic scales. Nevertheless, we note that even in full-PIC simulations some caution is warranted when interpreting the results, as residual dependencies on the adopted mass ratio may persist. This comparison, shown in Figure 1, serves to validate resolution convergence by benchmarking the hybrid approach against simulations that include the full electron kinetic physics. Figure 1 shows also the result obtained from a 2D simulation at $M_{A}=30$ with an intermediate resolution of $\Delta x=0.2$ $d_{i}$. The resolution of $\Delta x=0.1$ $d_{i}$ accurately captures the filaments produced by the ion-Weibel instability (E. S. Weibel, 1959; A. Bohdan et al., 2021; T. Jikei et al., 2024; L. Orusa & T. Jikei, 2026) that mediates the shock transition. These filaments are subsequently compressed and amplified by the shock, a process not properly captured at lower or intermediate resolutions.
@@
The choice of numerical resolution is a critical aspect of hybrid-PIC simulations. One may wonder whether further increasing the resolution would suppress particle acceleration or lead to more reliable results. This is not necessarily the case. The resolution adopted here ($\Delta x=0.1\,d_{i}$) was chosen because it reproduces the ion-scale magnetic-field structure observed in corresponding full-PIC simulations, which we use as a benchmark for physical fidelity. Increasing the resolution beyond this range does not automatically improve the accuracy of the hybrid approach. Instead, it can introduce numerical artifacts, most notably the onset of non-physical whistler modes due the fact that hybrid assumes massless $e^{-}$, which prevent convergence (L. Orusa & T. Jikei, 2026). This behavior limits the range of resolutions over which hybrid simulations provide a faithful description of the ion-scale instabilities relevant to quasi-perpendicular shocks.
Figure 3: Trajectories in the xyx-y plane (top) and xzx-z plane (bottom) of one representative particle per simulation—both 2D (dashed) and 3D (solid), at high and low resolution of $M_{A}=30$ shock—that exhibit similar behavior in the initial stages of their evolution (see the text for details).
III.2 The life of an accelerated particle in a perpendicular shock
Figure 4: The average penetration depths of ions with energy $E=7.5E_{sh}$ are shown for $M_{A}=30$ (red lines) and 100 (black lines) in units of $r_{L}^{*}=r_{L}(7.5E_{sh})$ (see the text for details). These represent the distances traveled downstream where $p_{x}$ changes sign. They are reported for 2D and 3D simulations at low and high resolutions. The hatched bands mark the $2\sigma$ intervals of the penetration distance distributions. The dotted line represents the 2D return horizon, as defined in Section III.3.
To investigate the differences in particle injection between 2D and 3D simulations at both low and high resolutions, we present in Figure 3 the xxyy (top) and xzx–z (bottom) trajectories of one representative particle from each simulation,comparing 2D and 3D, both at high and low resolution for $M_{A}=30$. These particles are selected because they exhibit similar early-time behavior, including comparable distances traveled ahead of the shock after the first reflection, similar energies after the first gyration, low $y$-momentum following the initial cycle, and similar maximum penetration depths into the downstream region. These “doppelgänger” particles serve as analogs across different simulation setups, enabling a direct comparison of their dynamics.
@@ III.2 The life of an accelerated particle in a perpendicular shock
Each particle undergoes a first reflection at the shock and propagates a similar distance upstream, between $17$ and 2020 $d_{i}$, gaining energy from the motional electric field. Following their initial reflection, they predominantly gyrate in the xzx–z plane due to the background magnetic field $\mathbf{B_{0}}=B_{0}\mathbf{\hat{y}}$, with a positive momentum $p_{z}$ and a small $p_{y}$, reaching energies around $7E_{sh}$ in this first SDA cycle. After this stage their evolutions begin to diverge. Once they penetrate in the downstream, all particles travel comparable distances before their $p_{x}$ switches from negative back to positive; we call this distance, which ranges between 20 and 25 $d_{i}$ downstream of the shock, the maximum penetration depth.
@@
III.3 The return horizon
To assess this porosity argument, we perform two experiments in which test particles (40000 particles in each test) are propagated through the electromagnetic fields extracted from the simulations at different times (8, 16, 24, and 32 $\omega_{c}^{-1}$) which yield consistent results with one another. The motion of the shock is accounted for by translating the shock position at each timestep, so that particles experience the correct relative dynamics. We use ions with energy $E=7.5\,E_{sh}$; particles with this energy have already undergone the first SDA cycle and above this threshold differences between 2D and 3D spectra become significant.
@@ III.3 The return horizon
We begin with the first test. Due to the upstream fields, reflected thermal particles predominantly gyrate in the xzx–z plane and enter the shock with large $p_{x,z}$ components and a smaller $p_{y}$ after the first SDA cycle. We initialize test particles at the shock and model their momentum by sampling $p_{x}^{2}$, $p_{y}^{2}$, and $p_{z}^{2}$ from uniform distributions. The sampling intervals are based on the phase spaces measured in the simulations; all test particles are constrained to have the same energy of $E=7.5\,E_{sh}$. Specifically, we use $p_{y}^{2}/p^{2}\in[0.0\,;0.2]$ with both positive and negative $p_{y}$, $p_{x}^{2}/p^{2}\in[0.0\,;0.4]$ with $p_{x}<0$ (since the particle is entering the shock), and $p_{z}>0$ obtained as $p_{z}=\sqrt{p^{2}-p_{x}^{2}-p_{y}^{2}}$.
@@
To identify this horizon, we consider particles initialized at their maximum penetration depth, where $p_{x}=0$, and assess how the return probability depends on the starting distance behind the shock. We explore two scenarios: one in which the particle momentum is aligned along the $z$-direction, and another in which particles have a finite spread in $p_{y}$. In the latter case, $p_{y}^{2}$ is sampled from a uniform distribution and normalized such that the total particle energy is set at $E=7.5,E_{sh}$, with $p_{y}^{2}/p^{2}=[0.0\;0.2]$ or $p_{y}^{2}/p^{2}=[0.0\;0.4]$ used as limiting values. We introduce a finite spread in the $y$-direction to demonstrate that, regardless of the initial momentum distribution, the return horizon remains fundamentally different between 2D and 3D.
@@
Figure 5 shows the percentage of particles with initial energy $E=7.5\,E_{sh}$ that, starting from various distances behind the shock, are able to return upstream. Results are presented for both 2D and 3D simulations at high (left panel) and low (right panel) resolution for $M_{A}=30$ and $M_{A}=100$. A clear contrast emerges: in 2D, the return probabilities rapidly drop to zero at the return horizon situated at $1.31\,r_{L}^{*}$ for both values of $M_{A}$, indicating a strong advective effect that prevents particles from escaping the downstream region. In 3D, however, a significant fraction of particles starting well beyond $1.31\,r_{L}^{*}$ are able to return upstream, consistent with the behavior observed in the “doppelgänger” particle trajectories of Section III.2. The 2D return horizon approximately coincides with the lower boundary of the $2\sigma$ band of the distribution of penetration depths, implying that most particles propagate downstream to distances beyond the return horizon, as shown in Figure 4. This explains the inability of particles to return upstream in 2D.
Figure 6: Spatial distribution (white contours) in the yyzz plane of test particles from our horizon exercise (based on the setup used for Figure 5) that successfully return upstream. The snapshot is taken at a position 0.3 $r_{L}^{*}$ behind the shock (with the test particles initially located at 1.6 $r_{L}^{*}$), for both high-resolution (left) and low-resolution (right) simulations at $M_{A}=30$. This distribution is compared to a snapshot of $B_{\perp}$ at the same location, illustrating that particles capable of returning upstream tend to avoid regions of strong $B_{\perp}$.
@@ III.4 The shock porosity
Figure 7: Filling factor $F(x,y,z)$ distribution obtained for 2D (top row) and 3D (bottom row) at $M_{A}=30$ high (left) and low resolution (right) as a function of the downstream distance from the shock. $F(x,y,z)$ is computed at each yyzz coordinate for different $x$ from the shock. From this sample, the whole distribution is obtained. A small value of $F$ implies the presence of regions of weak $B_{\perp}/B_{0}$. The comparison clearly shows that in 3D, $F$ tends to be systematically lower than in 2D, indicating enhanced porosity.
To support this argument, we consider test particles that successfully return upstream from an initial distance of 1.6 $r_{L}^{*}$; their spatial distribution in the yyzz plane at $x_{sh}-x=0.3\,r_{L}^{*}$ is shown in Figure 6 as white contours, together with the slice of $B_{\perp}$ taken at the same location, for both high- and low-resolution simulations with $M_{A}=30$. It is evident, especially for the low resolution simulation, that particles capable of returning upstream tend to avoid regions of strong $B_{\perp}$, instead favoring areas where $B_{\perp}$ is small.
To quantify the occurrence of such regions, we perform a statistical analysis of the filling factor of small-$B_{\perp}$ regions in the near downstream. In particular, we assess the shock porosity by searching for extended regions—“channels” or “tubes”—at fixed $y$ and $z$ coordinates where $B_{\perp}$ remains consistently below a defined threshold along the $x$ direction, all the way to the shock. The presence of such structures implies reduced magnetic confinement and enables particle propagation back towards the shock. Although particle trajectories are fully three-dimensional and not confined to fixed yyzz locations, this test provides both qualitative and quantitative evidence for the existence of preferential escape pathways.
@@
$f(x^{\prime},y,z)=\begin{cases}1&\text{if }B_{\perp}(x^{\prime},y,z)/B_{0}<T,\\ 0&\text{otherwise}.\end{cases}$ (2)
@@
In 2D, the translational invariance in the $z$ direction implies that the magnetic field is organized in ”walls”, with the magnetic structures forming extended, sheet-like barriers that span the simulation domain and systematically impede particle transport. In 3D, by contrast, the amplified magnetic field occupies a smaller volume fraction, leaving localized regions of weak field in which particles experience reduced magnetic deflection. In the low-resolution case, there are extended regions of very weak magnetic field, for example between $z\simeq 7.5$ and $15\,d_{i}$ in Fig. 8. In the high-resolution simulation, while localized regions reach field strengths of order $\sim 15\,B_{0}$, there are also intervals (e.g., between $z\simeq 7$ and $14\,d_{i}$ in Fig. 8) where the field is significantly weaker, including regions without strong amplification. As a result, particles in 3D can propagate over distances comparable to their Larmor radius without encountering strong magnetic barriers. In contrast, in 2D a magnetic barrier inevitably intersects particle trajectories at some point, limiting their motion. This behavior is consistent with the trajectories shown in Fig. 3 that are not affected by visible distortions of the orbits (F. C. Jones et al., 1998).